# Fundamental limitations on the recoverability of quantum processes

### v1

@@ Appendix E Proof of monotonicity of $\mathbf{D}^{(n)}$
We restrict ourselves to the class of $(n+1)^{\text{th}}$ order processes which satisfies the following equation.
@@
For this class of $(n+1)^{\text{th}}$ order process, our definition of generalized divergence of two $n^{\text{th}}$ order processes satisfies the monotonicity relation (87). The proof goes as follows. Consider two $n^{\text{th}}$ order processes $\Theta^{(n)}$ and $\Xi^{(n)}$ such that action of any $(n+1)^{\text{th}}$ order process, say $\Phi^{(n+1)}$, can be written in terms of concatenation of $n^{\text{th}}$ order process given by Eq. (158). Now, we show that in this scenario data processing inequality will hold, i.e., $\mathbf{D}^{(n)}\left[\Theta^{(n)}\;\middle\|\;\Gamma^{(n)}\right]\geq\mathbf{D}^{(n)}\left[\Phi^{(n+1)}\left(\Theta^{(n)}\right)\;\middle\|\;\Phi^{(n+1)}\left(\Gamma^{(n)}\right)\right]$. Using Eq. (86), we can write
@@
Using this inequality in (160) we finally have

### v2

@@ Appendix E Proof of monotonicity of $\mathbf{D}^{(n)}$
We restrict ourselves to the class of $(n+1)^{\text{th}}$ order processes which satisfy the following equation.
@@
For this class of $(n+1)^{\text{th}}$ order processes, our definition of generalized divergence between two $n^{\text{th}}$ order processes satisfies the monotonicity relation, Eq.(87). The proof goes as follows. Consider two $n^{\text{th}}$ order processes $\Theta^{(n)}$ and $\Gamma^{(n)}$ such that action of any $(n+1)^{\text{th}}$ order process, say $\Phi^{(n+1)}$, can be written in terms of concatenation of $n^{\text{th}}$ order processes, given by Eq. (159). Now, we show that in this scenario, data processing inequality holds, i.e., $\mathbf{D}^{(n)}\left[\Theta^{(n)}\;\middle\|\;\Gamma^{(n)}\right]\geq\mathbf{D}^{(n)}\left[\Phi^{(n+1)}\left(\Theta^{(n)}\right)\;\middle\|\;\Phi^{(n+1)}\left(\Gamma^{(n)}\right)\right]$. Using Eq. (86), we can write
@@
Using this inequality in (161), we finally have